OpenStudy (anonymous):

Find the solution -3x-2y=15 8x+6y=-34

OpenStudy (anonymous):

use substitution

OpenStudy (lυἶცἶ0210):

Or you could try elimination, which ever works for you. \(\large -3x-2y=15 \) \(\large 8x+6y=-34 \) --------------------------------------- \(\large (-3x-2y=15)(3) \) \(\large 8x+6y=-34 \) -------------------------------------- \(\large -9x-6y=45 \) \(\large 8x+6y=-34 \) Now you can solve for y

OpenStudy (anonymous):

I just can't seem to grasp an understanding I'm so confused

OpenStudy (anonymous):

I never understood the x y thing

OpenStudy (lυἶცἶ0210):

What exactly about it is confusing, solving it?